# Weighted wave envelope estimates for the parabola

### v1

@@ Abstract.
In this paper, we extend Fefferman’s classical square function estimate for the parabola to a weighted setting. Our weighted square function estimate is derived from a weighted wave envelope estimate for the parabola. The bounds are formulated in terms of families of multiscale tubes together with weight parameters that quantify the distribution of the weight. As an application, we obtain some weighted $L^{p}$-estimates for a class of Fourier multiplier operators and for solutions to free Schrödinger equation.
@@ 1. Introduction
This inequality relies on the geometric observation by Fefferman [14] that the algebraic differences $\theta-\theta^{\prime}$ overlap only finitely often as $\theta\neq\theta^{\prime}$ vary. See also [17, 30, 20] for extensions to non-degenerate curves in higher dimensions. Square function estimates of the form (1) have several important applications in harmonic analysis. The sharp square function estimate (1) is known to imply sharp results for the Kakeya maximal function, the Bochner-Riesz multipliers, the Fourier restriction operator, and local smoothing estimates for the Schrödinger equation; see [6, 41] and references therein. In higher dimensions, it is conjectured that (1) holds with $L^{4}(\mathbb{R}^{2})$ replaced by $L^{\frac{2d}{d-1}}(\mathbb{R}^{d})$, which remains wide open.
@@
Here we mean by $A\lessapprox B$ an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$. The term $R^{-100}$ on the right-hand side of (4) is negligible for interesting weights $H$. For instance, we have $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\gg R^{-100}$ whenever $H$ is the characteristic function of a union of unit balls. We also use the notation $A\lesssim B$ to denote $A\leq CB$ with an absolute constant $C>0$, possibly depending on parameters such as $p$ and $\alpha$.
@@
For $H\equiv 1$, we have $\kappa_{p,H}(U)=1$ for any $U\in\mathbb{U}$. Thus, when $p=4$, Theorem 1.1 essentially recovers the classical square function estimate (1). For $2\leq p<4$, the $H\equiv 1$ case of Theorem 1.1 recovers square function estimates due to Gan [15], where more general small cap square function estimates are established. Our weighted square function estimates are inspired by weighted decoupling inequalities for the paraboloids, which have been extensively studied in recent years and applied to problems such as the Falconer distance set conjecture and Bochner-Riesz means; see, e.g., [18, 11, 16, 26] and references therein.
@@
Their wave envelope estimate for $\Gamma$ implies, among other consequences, the sharp $L^{4}$ square function estimate for $\Gamma$ and the sharp local smoothing estimate for the wave equation in 2+1 dimension. The $p=4$ and $H\equiv 1$ case of Theorem 1.4 recovers the $L^{4}$ wave envelope estimate [19, Equation (8)], which is implicit in [21]. We note that [19] established more refined versions of the $L^{4}$ wave envelope estimates, termed amplitude-dependent wave envelope estimates, for both the parabola $\mathcal{P}$ and the cone $\Gamma$.
@@
$\bullet$ We write $A\lessapprox B$ to denote an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$ for $R>1$ where $C_{p}$ is a constant depending only on $p$.
@@ 2.3. Weighted estimates for the Schrödinger equation
A detailed proof of (14) can be found in a note by Yung [41, Proof of Theorem 2]. It follows from a standard duality argument similar to the one used in [9] and [33]. For completeness, we sketch the argument in Section 5.3. ∎
@@
[…]
@@ 2.5. Fractal local smoothing estimates
In [28, Theorem 1.6], it is shown that if $\gamma\geq-\frac{\alpha}{2}$, then $H^{\gamma}$$L_{t}^{2}(d\nu;L_{x}^{2}(-1,1))$ estimates holds for $e^{it\partial_{x}^{2}}f$. Moreover, optimal results in higher dimensions and weighted Strichartz-type estimates of the form $H^{\gamma}$$L_{t}^{q}(d\nu;L_{x}^{r})$ for fractional Schrödinger operators are established in [28] (see also [4] for analogous results for the wave operator).
@@ 4.2. Decomposition
For $R\geq 1$, let $K$ be a dyadic number such that $K\sim\log R$. We choose $m\in\mathbb{N}$ such that $K^{m}\sim R^{1/2}$. We consider larger canonical covering $\tau$ such that $|\tau|=s$ for each scale
@@
At the smallest scale for $j=m$, we obtain the collection of blocks $\{\theta\}$ of size $|\theta|=K^{-m}\sim R^{-1/2}.$ For each block $\tau$ at an intermediate scale, we define
@@
We use the following pointwise estimate from [10, Section 5].
@@
For any $x\in\mathbb{R}^{2}$, there is an absolute constant $C>0$ (independent of $x$) such that
(32) $|f(x)|^{p}\leq C^{m}\sum_{\theta}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}|f_{\tau_{1}}(x)f_{\tau_{2}}(x)|^{\frac{p}{2}},$
@@
(33) $\|f\|_{L^{p}(Y)}^{p}\!\leq\!C^{m}\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}\!\!\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p}.$
@@
Since for each fixed $\tau$ there are only $K^{O(1)}$ (and hence $\lesssim\log R$) such pairs $\tau_{1},\tau_{2}\subset\tau$, this logarithmic factor can be absorbed into the $O((\log R)^{O(1)})$ loss. Consequently, the bilinear term in the sum in (33) is bounded by the right-hand side of (8).
@@ 4.4. Bilinear restriction estimates
The proof of Theorem 4.4 is standard; see, for example, [38] or [29, Lemma 2.4].
@@ 6. Examples for lower bounds
Let $R\geq 1$. Recall that $K$ is a dyadic number chosen so that
$1\leq K\leq\dots\leq K^{m}\sim R^{1/2}.$
@@
$|f(x)|^{p}\leq C\max_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}(x)|^{p}+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}$
@@
$|f(x)|^{p}$ $\leq C^{2}\max_{\tau_{1}}\max_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}|f_{\tau_{2}}(x)|^{p}+C^{2}K^{p}\max_{\tau_{1}}\max_{\begin{subarray}{c}\tau_{2},\tau_{2}^{\prime}\in\mathcal{T}_{2}(\tau_{1});\\ d(\tau_{2},\tau_{2}^{\prime})\geq\frac{1}{2K^{2}}\end{subarray}}|f_{\tau_{2}}(x)|^{\frac{p}{2}}|f_{\tau_{2}^{\prime}}(x)|^{\frac{p}{2}}$
$\qquad\quad+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}.$
Continuing in this manner, with $\mathcal{T}:=\mathcal{T}_{m}$, we get
$\begin{split}|f(x)|^{p}&\leq C^{m}\max_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\max_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}\\ &\leq C^{m}\sum_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\sum_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}.\end{split}$
This completes the proof of Lemma 4.1. ∎
@@ Acknowledgement
J. Kim was supported in part by grants from the Research Grants Council of the Hong Kong Administrative Region, China (Project No. CityU 21309222 and CityU 11308924). H. Ko was supported in part by the NRF (Republic of Korea) grant RS-2024-00339824, G-LAMP RS-2024-00443714, JBNU research funds for newly appointed professors in 2024, and NRF2022R1I1A1A01055527. We thank the organizers of the Special Topic School “Maximal Operators and Applications” and the Hausdorff Center for Mathematics, where our collaboration began. We also thank Sanghyuk Lee for helpful discussions. The second author would like to thank Seheon Ham for previous collaboration related to this topic.

### v2

@@ Abstract.
In this paper, we extend the Córdoba–Fefferman square function estimate for the parabola to a weighted setting. Our weighted square function estimate is derived from a weighted wave envelope estimate for the parabola. The bounds are formulated in terms of families of multiscale tubes together with weight parameters that quantify the distribution of the weight. As an application, we obtain some weighted $L^{p}$-estimates for a class of Fourier multiplier operators and for solutions to free Schrödinger equation.
@@ 1. Introduction
This inequality relies on the geometric observation by Fefferman [16] that the algebraic sums $\theta+\theta^{\prime}$ overlap only finitely often as $\theta$ and $\theta^{\prime}$ vary. See also [19, 32, 22] for extensions to non-degenerate curves in higher dimensions. Square function estimates of the form (1) have several important applications in harmonic analysis. The sharp square function estimate (1) is known to imply sharp results for the Kakeya maximal function, the Bochner-Riesz multipliers, the Fourier restriction operator, and local smoothing estimates for the Schrödinger equation; see [6, 43] and references therein. For the paraboloid in higher dimensions, it is conjectured that (1) holds with $L^{4}(\mathbb{R}^{2})$ replaced by $L^{\frac{2d}{d-1}}(\mathbb{R}^{d})$, which remains wide open.
@@
Here we mean by $A\lessapprox B$ an inequality of the form $A\leq C_{\epsilon}R^{\epsilon}B$ for any $\epsilon>0$. The term $R^{-100}$ on the right-hand side of (4) is negligible for interesting weights $H$. For instance, we have $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\gg R^{-100}$ whenever $H$ is the characteristic function of a union of unit balls. We also use the notation $A\lesssim B$ to denote $A\leq CB$ with an absolute constant $C>0$, possibly depending on parameters such as $p$ and $\alpha$.
@@
For $H\equiv 1$, we have $\kappa_{p,H}(U)=1$ for any $U\in\mathbb{U}$. Thus, when $p=4$, Theorem 1.1 essentially recovers the classical square function estimate (1). For $2\leq p<4$, the $H\equiv 1$ case of Theorem 1.1 recovers square function estimates due to Gan [17], where more general small cap square function estimates are established. The case $p=2$ of Theorem 1.1 is essentially contained in [8]; this was kindly pointed out to us by Tony Carbery (see also [7, pages 17–19]). Our weighted square function estimates are inspired by weighted decoupling inequalities for the paraboloids, which have been extensively studied in recent years and applied to problems such as the Falconer distance set conjecture and Bochner-Riesz means; see, e.g., [20, 13, 18, 28] and references therein.
@@
Their wave envelope estimate for $\Gamma$ implies, among other consequences, the sharp $L^{4}$ square function estimate for $\Gamma$ and the sharp local smoothing estimate for the wave equation in 2+1 dimension. The $p=4$ and $H\equiv 1$ case of Theorem 1.4 recovers the $L^{4}$ wave envelope estimate [21, Equation (8)], which is implicit in [23]. We note that [21] established more refined versions of the $L^{4}$ wave envelope estimates, termed amplitude-dependent wave envelope estimates, for both the parabola $\mathcal{P}$ and the cone $\Gamma$.
@@
$\bullet$ We write $A\lessapprox B$ to denote an inequality of the form $A\leq C_{\epsilon}R^{\epsilon}B$ for any $\epsilon>0$ where $C_{\epsilon}$ is an absolute constant.
@@ 2.3. Weighted estimates for the Schrödinger equation
A detailed proof of (14) can be found in a note by Yung [43, Proof of Theorem 2]. It follows from a standard duality argument similar to the one used in [11] and [35]. For completeness, we sketch the argument in Section 5.3. ∎
@@
2.3 can be obtained by interpolation between known $L^{2}$ and $L^{4}$ estimates. Indeed, when $p=2$, 2.3 recovers a bound due to Du and Zhang [14] (see also [42, 15]). The $p=4$ case of 2.3 can be deduced from a local smoothing estimate for the Schrödinger equation:
@@ 2.5. Fractal local smoothing estimates
In [30, Theorem 1.6], it is shown that if $\gamma\geq-\frac{\alpha}{2}$, then $H^{\gamma}$$L_{t}^{2}(d\nu;L_{x}^{2}(-1,1))$ estimates holds for $e^{it\partial_{x}^{2}}f$. Moreover, optimal results in higher dimensions and weighted Strichartz-type estimates of the form $H^{\gamma}$$L_{t}^{q}(d\nu;L_{x}^{r})$ for fractional Schrödinger operators are established in [30] (see also [4] for analogous results for the wave operator).
@@ 4.2. Decomposition
Let $K\sim\log R$ be a dyadic number. Let $m\in\mathbb{N}$ be the integer such that $K^{m}\leq R^{1/2}<K^{m+1}$. We consider a canonical covering $\{\tau\}$ such that $|\tau|=s$ for each scale
@@
For each block $\tau$ at an intermediate scale, we define
@@
We use the following pointwise estimate from [12, Section 5].
@@
For any $x\in\mathbb{R}^{2}$, we have the pointwise bound
(32) $|f(x)|^{p}\lessapprox\sum_{|\theta|=R^{-1/2}}|f_{\theta}(x)|^{p}+\sum_{1\leq j\leq m}\sum_{|\tau|=K^{-j+1}}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-j}\\ d(\tau_{1},\tau_{2})\geq K^{-j}\end{subarray}}|f_{\tau_{1}}(x)f_{\tau_{2}}(x)|^{\frac{p}{2}},$
@@
(33) $\|f\|_{L^{p}(Y)}^{p}\lessapprox\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}+\sum_{1\leq j\leq m}\sum_{|\tau|=K^{-j+1}}\max_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-j}\\ d(\tau_{1},\tau_{2})\geq K^{-j}\end{subarray}}\!\!\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p},$
where we have used the fact that, for each fixed $\tau$, there are at most $K^{O(1)}\lessapprox 1$ pairs $\tau_{1},\tau_{2}\subset\tau$.
@@
This estimate bounds the bilinear terms on the right-hand side of (33) by the right-hand side of (8).
@@ 4.4. Bilinear restriction estimates
The proof of Theorem 4.4 is standard; see, for example, [40] or [31, Lemma 2.4].
@@ 6. Examples for lower bounds
Recall that $K\sim\log R$ is a dyadic number and $m$ is the integer such that
$1\leq K\leq\dots\leq K^{m}\leq R^{1/2}<K^{m+1}.$
@@
$|f(x)|^{p}\leq C\max_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}(x)|^{p}+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq K^{-1}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}$
@@
$|f(x)|^{p}$ $\leq C^{2}\max_{\tau_{1}}\max_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}|f_{\tau_{2}}(x)|^{p}+C^{2}K^{p}\max_{\tau_{1}}\max_{\begin{subarray}{c}\tau_{2},\tau_{2}^{\prime}\in\mathcal{T}_{2}(\tau_{1});\\ d(\tau_{2},\tau_{2}^{\prime})\geq K^{-2}\end{subarray}}|f_{\tau_{2}}(x)|^{\frac{p}{2}}|f_{\tau_{2}^{\prime}}(x)|^{\frac{p}{2}}$
$\qquad\quad+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq K^{-1}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}.$
Continuing in this manner, we get
$\begin{split}|f(x)|^{p}&\leq C^{m}\max_{\tau_{m}\in\mathcal{T}_{m}}|f_{\tau_{m}}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\max_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq K^{-j}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}\\ &\lessapprox\sum_{|\theta|=R^{-1/2}}|f_{\theta}(x)|^{p}+\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\sum_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq K^{-j}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}},\end{split}$
where we have used $C^{m}K^{O(1)}\lessapprox 1$ together with the fact that each $\tau_{m}\in\mathcal{T}_{m}$ contains $O(K)$ elements $\theta$ of size $|\theta|=R^{-1/2}$. This completes the proof of Lemma 4.1. ∎
@@ Acknowledgement
J. Kim was supported in part by grants from the Research Grants Council of the Hong Kong Administrative Region, China (Project No. CityU 21309222 and CityU 11308924). H. Ko was supported in part by the NRF (Republic of Korea) grant RS-2024-00339824, G-LAMP RS-2024-00443714, JBNU research funds for newly appointed professors in 2024, and NRF2022R1I1A1A01055527. We thank the organizers of the Special Topic School “Maximal Operators and Applications” and the Hausdorff Center for Mathematics, where our collaboration began. We also thank Sanghyuk Lee for helpful discussions. The second author would like to thank Seheon Ham for previous collaboration related to this topic. We are grateful to Tony Carbery for bringing the references [8, 7] to our attention and for pointing out the overlap between their results and the case $p=2$ of Theorem 1.1.